# Amortized Safe Active Learning for Real-Time Decision-Making: Pretrained Neural Policies from Simulated Nonparametric Functions

### v1

@@ 3.1.1 Unconstrained AL: Simulated Acquisition
Assume for now that safety conditions are not considered. We are then dealing with an unconstrained AL problem: we collect $y_{\phi,1:T_{\text{sim}}}$ informative for $f$. The GP prior $f\sim\mathcal{GP}(\bm{0},k_{\theta})$ further turns this into an active GP learning problem. This means common acquisition functions (Seo et al., 2000; Guestrin et al., 2005; Krause et al., 2008) are valid to guide our queries to high information. Note, however, that we now optimize w.r.t. the policy where gradient is propagated from all queries jointly. A good choice here is the entropy (Seo et al., 2000; Krause & Guestrin, 2007), which (i) has closed forms, (ii) is differentiable, and (iii) is one of the gold standard acquisition criteria for GPs.
@@
where $\mathbb{E}_{f,T_{\text{sim}},\mathcal{E}_{\text{init}},\epsilon_{1:T_{\text{sim}}}}\left[h(y_{\phi,1:T_{\text{sim}}},Y_{\text{init}})\right]$, an average of various instances, represents the policy’s entropy defined by Krause & Guestrin (2007) (originally for conventional AL). $p(\cdot)$ is a GP likelihood (Appendix A). The proportion symbol here indicates equivalency which holds by applying Bayes rule and removing the part that has no gradient w.r.t. the policy. The effect of $T_{\text{sim}}$ will be described shortly after, let us say $T_{\text{sim}}=T$ is fixed here. Maximizing this acquisition criterion means the policy selects points that are the most distinctive to each other. We will later see that this choice has the advantage of explainability in conjunction with safety constraints. In our Appendix C (Figure A.C.5), we illustrate the acquisition value with $T=1$.
The ideas until here are similar to Chen et al. (2017); Foster et al. (2021); Ivanova et al. (2021), i.e. turning the acquisition criteria we would have optimized sequentially into trainable objectives in an a priori simulated learning.
@@
$\begin{split}I(y_{\phi,1:T_{\text{sim}}}|Y_{\text{init}})=&-\log p(y_{\phi,1:T_{\text{sim}}}|Y_{\text{init}})\\ &+\log p(y_{\phi,1:T_{\text{sim}}}|Y_{\text{init}},Y_{\text{grid}}),\end{split}$ (3)
where $Y_{\text{grid}}$ are noisy evaluations at randomly sampled $\bm{X}_{\text{grid}}\subseteq\mathcal{X}$ (with $|\bm{X}_{\text{grid}}|\gg T$). This $I(\cdot)$ is adapted from the mutual information AL criterion (Guestrin et al., 2005; Krause et al., 2008). It encourages the policy to look inside the input space and avoids over-emphasizing the border of $\mathcal{X}$, a problem of entropy. We put details into Appendix C.
@@ 3.1.2 Unconstrained AL Objective, beyond Fixed Length & Fixed Priors
One remark of the previous objectives is that they are nonmyopic [^1] Nonmyopic exploration allocates multiple points jointly, e.g. query $1/3,2/3$ in $[0,1]$ instead of greedily selecting $1/2,1/4$. (Krause & Guestrin, 2007; Foster et al., 2021), assuming a pre-defined budget $T$. The resulting queries can be suboptimal if we deploy for fewer steps. In practice, it is often unrealistic to know the precise number of queries $T$ in advance, especially if we wish to deploy the trained policy on multiple problems. To this end, we assign random AL budget to each $f$, i.e. $T_{\text{sim}}\sim\text{Uniform}[1,T]$. Then, the exploration scores are normalized by the sequence length:
$\mathcal{H}(\phi)=\mathbb{E}_{\theta,\sigma^{2}}$ $\mathbb{E}_{f,T_{\text{sim}},\mathcal{E}_{\text{init}},\epsilon_{1:T_{\text{sim}}}}\left[\frac{h(y_{\phi,1:T_{\text{sim}}}|Y_{\text{init}})}{N_{\text{init}}+T_{\text{sim}}}\right],$ (4)
$\mathcal{I}(\phi)=\mathbb{E}_{\theta,\sigma^{2}}$ $\mathbb{E}_{f,T_{\text{sim}},\mathcal{E}_{\text{init}},\epsilon_{1:T_{\text{sim}}}}\left[\frac{I(y_{\phi,1:T_{\text{sim}}|Y_{\text{init}})}}{N_{\text{init}}+T_{\text{sim}}}\right].$ (5)
@@
0:  2.1, $T$, $N_{\text{init}}$
1:  sample a batch of $(\theta,\sigma^{2},\theta_{q},\sigma_{q}^{2})$
2:  sample a batch of $(f,q)$ pairs per Algorithm 3
3:  for each $f,q$ do
4:     given $\mathcal{D}_{0}$ per Algorithm 3
5:     sample $T_{\text{sim}}\sim\text{Uniform}[1,T]$
6:     for $t=1,...,T_{\text{sim}}$ do
7:        $\bm{x}_{\phi,t}=\phi(T_{\text{sim}}-t+1,\mathcal{D}_{t-1})$
8:        sample $\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2}),y_{\phi,t}=f(\bm{x}_{\phi,t})+\epsilon_{t}$
9:        sample $\epsilon_{q,t}\sim\mathcal{N}(0,\sigma_{q}^{2}),z_{\phi,t}=q(\bm{x}_{\phi,t})+\epsilon_{q,t}$
10:        $\mathcal{D}_{t}\leftarrow\mathcal{D}_{t-1}\cup\{\bm{x}_{\phi,t},y_{\phi,t},z_{\phi,t}\}$
11:     end for
@@
13:  compute loss per Eq. 7, update $\phi$
3.1.3 Safe AL Objective
We have obtained an entropy objective for unconstrained AL. Now we take $z_{\phi,1:T_{\text{sim}}}$ into consideration and we wish to follow the same intuition to get a safe AL objective. Recall that a conventional safe AL (Schreiter et al., 2015; Zimmer et al., 2018) solves Eq. 1 for each query. This constrained acquisition criterion has proven its effectiveness, and we next translate it into a differentiable objective compatible to our simulated environment.
@@ 3.1.3 Safe AL Objective
    The constraint $p(z(\bm{x})\geq 0|z_{\phi,1:t},Z_{\text{init}})\geq 1-\gamma$ is equivalent to $\log p(z(\bm{x})<0|z_{\phi,1:t},Z_{\text{init}})\leq\log(\gamma)$.
@@
[…]
    $-\lambda$ $\log p(z(\bm{x})<0|z_{\phi,1:t},Z_{\text{init}})+\lambda\log\gamma\},$
@@
    Since $\gamma$ had no effect in the original formulation ($\log\gamma$ is constant w.r.t. $\bm{x}$), we twist the dual problem into
[…]
    $-\log$ $\text{max}(\gamma,p(z(\bm{x})<0|z_{\phi,1:t},Z_{\text{init}}))\},$
@@
$\mathcal{S}_{\mathcal{H}}(\phi)$ $=\mathbb{E}_{\begin{subarray}{c}\theta,\sigma^{2},\\ \theta_{q},\sigma_{q}^{2}\end{subarray}}\mathbb{E}_{\begin{subarray}{c}f,T_{\text{sim}},\mathcal{E}_{\text{init}},\epsilon_{1:T_{\text{sim}}},\\ q,\mathcal{E}_{q,\text{init}},\epsilon_{q,1:T_{\text{sim}}}\end{subarray}}\left[\frac{S(\mathcal{D}_{T_{\text{sim}}})}{N_{\text{init}}+T_{\text{sim}}}\right],\text{with }$ (7)
$S(\mathcal{D}_{T_{\text{sim}}})$ $=-\log p(y_{\phi,1:T_{\text{sim}}}|Y_{\text{init}})$
$-\sum_{t=0}^{T_{\text{sim}}-1}\log$ $\text{max}(\gamma,p(z(\bm{x}_{\phi,t+1})<0|z_{\phi,1:t},Z_{\text{init}})).$
@@
The objective function $\mathcal{S}_{\mathcal{H}}(\phi)$ can vary with the order of queried data, as the order directly impacts the safety score. Note that the Eq. 6 can itself be a constraint-aware yet differentiable safe AL acquisition criterion for conventional setup (see ablation studies in Appendix G).
@@ 3.2 Function Sampling
0:  2.1, center $\mathcal{C}\subseteq\mathcal{X}$, $N_{\text{init}}$, $\mathcal{D}=\emptyset$, $\mathcal{GP}(0,k_{\theta}),\mathcal{GP}_{\theta_{q}}(\mu_{q},k_{q}),\sigma^{2},\sigma_{q}^{2}$, max_iter$=50$
1:  sample $f_{0}\sim\mathcal{GP}(0,k_{\theta})$, $q_{0}\sim\mathcal{GP}(0,k_{q})$
2:  $f(\cdot)=f_{0}(\cdot)-\mathbb{E}_{\bm{x}\in\mathcal{X}}[f_{0}(\bm{x})]$
3:  $q(\cdot)=\mu_{q}(\cdot)+q_{0}(\cdot)-\mathbb{E}_{\bm{x}\in\mathcal{X}}[q_{0}(\bm{x})]$
4:  for i=1,…, max_iter do
5:     sample $\bm{X}\sim\text{Uniform}[\mathcal{C}],|\bm{X}|=N_{\text{init}}$
6:     sample $\mathcal{E}_{\text{init}}\sim\mathcal{N}(0,\sigma^{2}I),Y=f(\bm{X})+\mathcal{E}_{\text{init}}$
7:     sample $\mathcal{E}_{q,\text{init}}\sim\mathcal{N}(0,\sigma_{q}^{2}I),Z=q(\bm{X})+\mathcal{E}_{q,\text{init}}$
8:     take those safe: $\mathcal{D}\leftarrow\mathcal{D}\cup[\bm{X},Y,Z]_{z\geq 0}$
9:     if i$=$max_iter then $\mathcal{D}\leftarrow\mathcal{D}\cup(\bm{X},Y,Z)$
10:     if $|\mathcal{D}|\geq N_{\text{init}}$ then take first $N_{\text{init}}$ and break
11:  end for
12:  return $f,q,\mathcal{D}_{0}=\mathcal{D}$ ($|\mathcal{D}|=N_{\text{init}}$)
@@ 3.2.1 Safety Function $q$ and Initial Data $\mathcal{D}_{0}$
The first challenge is to ensure that that the simulated $q$ and $\mathcal{D}_{0}$ are both realistic and representative. In safe exploration problems (Sui et al., 2015; Zimmer et al., 2018; Bottero et al., 2022), the initial data are usually provided by a domain expert at the centric area of a safe set and the algorithms gradually explore towards the safe set border. Following this principle, our generative algorithm samples initial data $\mathcal{D}_{0}$ in a pre-defined safe set, and we assume w.l.o.g. that the safe set lies at the center of $\mathcal{X}$, denoted by $\mathcal{C}\subseteq\mathcal{X}$. This approach ensures that $\mathcal{D}_{0}$ closely resembles the initial datasets encountered in the deployment stage.
@@ 3.2.2 Fourier Feature Functions - Efficient and Decoupled
The final step is to devise an efficient sampling strategy for the noisy GP values $y_{\phi,1:T_{\text{sim}}}$ and $z_{\phi,1:T_{\text{sim}}}$. This is not trivial because the observations are sampled iteratively, e.g. $\bm{x}_{t}=\phi(\mathcal{D}_{t-1})$, meaning $y_{\phi,1:t-1}$, $z_{\phi,1:t-1}$ are sampled before we know $\bm{x}_{t},...,\bm{x}_{T}$. One can make standard GP posterior sampling conditioned on preceding samples, e.g. $y_{t}\sim p(y(\bm{x}_{t})|\mathcal{D}_{t-1},k_{\theta},\sigma^{2})$. However, when performed naively, this posterior sampling approach is bound to the GP cubic complexity for each data point acquisition $y_{t}$, which is prohibitive. The runtime can be further reduced by using low-rank updates (Seeger, 2004). For specific objectives such as for our $\mathcal{H}$ or $\mathcal{S}_{\mathcal{H}}$, the GP posteriors can be then computed as intermediate results with minimal additional cost (see Eqs. 6 and 7). However, low-rank updates can lead to additional issues, as they make the computational graph unnecessarily complex for the loss function differentiation. To overcome these challenges, we propose a decoupled approach, which is scalable by itself and enables flexible integration of any objectives, e.g. potentially cheap approximated objectives, for future work.
The core is a decoupled function sampling technique (Rahimi & Recht, 2007; Wilson et al., 2020), where each GP function is approximated by a linear combination of Fourier features. The function value can be evaluated later at any $\bm{x}\in\mathcal{X}$ in linear time (line 8-9 of Algorithm 2). One limitation, however, is that the kernels $k_{\theta},k_{q}$ need to have Fourier transforms (e.g. stationary kernels, see Bochner’s theorem in Rasmussen & Williams, 2006). For $q\sim\mathcal{GP}(\mu_{q},k_{q})$, one may sample $q_{0}\sim\mathcal{GP}(0,k_{q})$ and set $q=\mu_{q}+q_{0}$. Note that a sampled function does not necessarily average to zero on $\mathcal{X}$, but we may compute a domain specific average analytically with usually negligible complexity (details in Section D.1). This is preferred because we consider normalized deployment problems, and such an optional mean shift improves the performance.

### v2

@@ 3.1.1 Unconstrained AL: simulated acquisition
Assume for now that safety conditions are not considered. We are then dealing with an unconstrained AL problem: we collect $y_{\phi,1:T_{\text{sim}}}$ informative for $f$. The GP prior $f\sim\mathcal{GP}(\bm{0},k_{\theta})$ further turns this into an active GP learning problem. This means common acquisition functions [42, 17, 23] are valid to guide our queries to high information. Note, however, that we now optimize w.r.t. the policy where gradient is propagated from all queries jointly. A good choice here is the entropy [42, 22], which (i) has closed forms, (ii) is differentiable, and (iii) is one of the gold standard acquisition criteria for GPs.
@@
where $\mathbb{E}_{f,T_{\text{sim}},\mathcal{E}_{\text{init}},\epsilon_{1:T_{\text{sim}}}}\left[h(y_{\phi,1:T_{\text{sim}}},Y_{\text{init}})\right]$, an average of various instances, represents the policy’s entropy defined by Krause & Guestrin [22] (originally for conventional AL). $p(\cdot)$ is a GP likelihood (Appendix A). The proportion symbol here indicates equivalency which holds by applying Bayes rule and removing the part that has no gradient w.r.t. the policy. The effect of $T_{\text{sim}}$ will be described shortly after, let us say $T_{\text{sim}}=T$ is fixed here. Maximizing this acquisition criterion means the policy selects points that are the most distinctive to each other. We will later see that this choice has the advantage of explainability in conjunction with safety constraints. In our Appendix C (Figure A.C.5), we illustrate the acquisition value with $T=1$.
The ideas until here are similar to Chen et al. [10], Foster et al. [11], Ivanova et al. [20], i.e. turning the acquisition criteria we would have optimized sequentially into trainable objectives in an a priori simulated learning.
@@
$\begin{split}I(y_{\phi,1:T_{\text{sim}}}|Y_{\text{init}})=-\log p(y_{\phi,1:T_{\text{sim}}}|Y_{\text{init}})+\log p(y_{\phi,1:T_{\text{sim}}}|Y_{\text{init}},Y_{\text{grid}}),\end{split}$ (3)
where $Y_{\text{grid}}$ are noisy evaluations at randomly sampled $\bm{X}_{\text{grid}}\subseteq\mathcal{X}$ (with $|\bm{X}_{\text{grid}}|\gg T$). This $I(\cdot)$ is adapted from the mutual information AL criterion [17, 23]. It encourages the policy to look inside the input space and avoids over-emphasizing the border of $\mathcal{X}$, a problem of entropy. We put details into Appendix C.
@@ 3.1.2 Unconstrained AL objective, beyond fixed length & fixed priors
One remark of the previous objectives is that they are nonmyopic [^1] Nonmyopic exploration allocates multiple points jointly, e.g. query $1/3,2/3$ in $[0,1]$ instead of $1/2,1/4$. [22, 11], assuming a pre-defined budget $T$. The resulting queries can be suboptimal if we deploy for fewer steps. In practice, it is often unrealistic to know the precise number of queries $T$ in advance, especially if we wish to deploy the trained policy on multiple problems. To this end, we assign random AL budget to each $f$, i.e. $T_{\text{sim}}\sim\text{Uniform}[1,T]$. Then, the exploration scores are normalized by the sequence length (AL sims indicate random $f,T_{\text{sim}},\mathcal{E}_{\text{init}},\epsilon_{1:T_{\text{sim}}}$):
$\mathcal{H}(\phi)=\mathbb{E}_{\theta,\sigma^{2}}\mathbb{E}_{\text{AL sims}}\left[\frac{h(y_{\phi,1:T_{\text{sim}}}|Y_{\text{init}})}{N_{\text{init}}+T_{\text{sim}}}\right],$ (4) $\mathcal{I}(\phi)=\mathbb{E}_{\theta,\sigma^{2}}\mathbb{E}_{\text{AL sims}}\left[\frac{I(y_{\phi,1:T_{\text{sim}}|Y_{\text{init}})}}{N_{\text{init}}+T_{\text{sim}}}\right].$ (5)
@@ 3.1.3 Safe AL objective
1:2.1, $T$, $N_{\text{init}}$
2:draw a batch of $(\theta,\sigma^{2},\theta_{q},\sigma_{q}^{2})$
3:draw a batch of $(f,q)$ pairs (Algorithm 3)
4:for each $f,q$ do
5:     given $\mathcal{D}_{0}$ per Algorithm 3
6:     draw $T_{\text{sim}}\sim\text{Uniform}[1,T]$
7:     for $t=1,...,T_{\text{sim}}$ do
8:         $\bm{x}_{\phi,t}=\phi(T_{\text{sim}}-t+1,\mathcal{D}_{t-1})$
9:         draw $\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2})$, $y_{\phi,t}=f(\bm{x}_{\phi,t})+\epsilon_{t}$
10:         draw $\epsilon_{q,t}\sim\mathcal{N}(0,\sigma_{q}^{2})$, $z_{\phi,t}=q(\bm{x}_{\phi,t})+\epsilon_{q,t}$
11:         $\mathcal{D}_{t}\leftarrow\mathcal{D}_{t-1}\cup\{\bm{x}_{\phi,t},y_{\phi,t},z_{\phi,t}\}$
@@
13:end for
14:compute loss per Eq. 8, update $\phi$
We have obtained an entropy objective for unconstrained AL. Now we take $z_{\phi,1:T_{\text{sim}}}$ into consideration and we wish to follow the same intuition to get a safe AL objective. Recall that a conventional safe AL [40, 55] solves Eq. 1 for each query. This constrained acquisition criterion has proven its effectiveness, and we next translate it into a differentiable objective compatible to our simulated environment.
@@
    $p(z(\bm{x})\geq 0|z_{\phi,1:t},Z_{\text{init}})\geq 1-\gamma$ is equivalent to $\log p(z(\bm{x})<0|z_{\phi,1:t},Z_{\text{init}})\leq\log(\gamma)$.
@@
    $\text{argmax}_{\bm{x}}\{-\log p(\hat{y}(\bm{x})|y_{\phi,1:t},Y_{\text{init}})-\lambda\log p(z(\bm{x})<0|z_{\phi,1:t},Z_{\text{init}})+\lambda\log\gamma\},$ (6)
@@
    Since $\gamma$ had no effect in Eq. 6 ($\log\gamma$ is constant w.r.t. $\bm{x}$), we twist the dual problem into
    $\text{argmax}_{\bm{x}}\{-\log p(\hat{y}(\bm{x})|y_{\phi,1:t},Y_{\text{init}})-\log\ \text{max}(\gamma,p(z(\bm{x})<0|z_{\phi,1:t},Z_{\text{init}}))\},$ (7)
@@
$\mathcal{S}_{\mathcal{H}}(\phi)$ $=\mathbb{E}_{\theta,\sigma^{2},\theta_{q},\sigma_{q}^{2}}\mathbb{E}_{f,T_{\text{sim}},\mathcal{E}_{\text{init}},\epsilon_{1:T_{\text{sim}}},q,\mathcal{E}_{q,\text{init}},\epsilon_{q,1:T_{\text{sim}}}}\left[\frac{S(\mathcal{D}_{T_{\text{sim}}})}{N_{\text{init}}+T_{\text{sim}}}\right],\text{with }$ (8)
$S(\mathcal{D}_{T_{\text{sim}}})$ $=-\log p(y_{\phi,1:T_{\text{sim}}}|Y_{\text{init}})-\sum_{t=0}^{T_{\text{sim}}-1}\log\ \text{max}(\gamma,p(z(\bm{x}_{\phi,t+1})<0|z_{\phi,1:t},Z_{\text{init}})).$
@@
The objective function $\mathcal{S}_{\mathcal{H}}(\phi)$ can vary with the order of queried data, as the order directly impacts the safety score. Note that the Eq. 7 can itself be a safety-aware yet unconstrained differentiable acquisition criterion for conventional safe AL setup (see ablation studies in Section G.4).
@@ 3.2 Function sampling
1:2.1, center $\mathcal{C}\subseteq\mathcal{X}$, $N_{\text{init}}$, $\mathcal{D}=\emptyset$, $\mathcal{GP}(0,k_{\theta}),\mathcal{GP}_{\theta_{q}}(\mu_{q},k_{q}),\sigma^{2},\sigma_{q}^{2}$, max_iter$=50$
2:draw $f_{\text{raw}}\sim\mathcal{GP}(0,k_{\theta})$, $q_{\text{raw}}\sim\mathcal{GP}(0,k_{q})$
3:$f(\cdot)=f_{\text{raw}}(\cdot)-\mathbb{E}_{\bm{x}\in\mathcal{X}}[f_{\text{raw}}(\bm{x})]$
4:$q(\cdot)=\mu_{q}(\cdot)+q_{\text{raw}}(\cdot)-\mathbb{E}_{\bm{x}\in\mathcal{X}}[q_{\text{raw}}(\bm{x})]$
5:for i=1,…, max_iter do
6:     draw $\bm{X}\sim\text{Uniform}[\mathcal{C}],|\bm{X}|=N_{\text{init}}$
7:     draw $\mathcal{E}_{\text{init}}\sim\mathcal{N}(0,\sigma^{2}I),Y=f(\bm{X})+\mathcal{E}_{\text{init}}$
8:     draw $\mathcal{E}_{q,\text{init}}\sim\mathcal{N}(0,\sigma_{q}^{2}I),Z=q(\bm{X})+\mathcal{E}_{q,\text{init}}$
9:     take those safe: $\mathcal{D}\leftarrow\mathcal{D}\cup[\bm{X},Y,Z]_{z\geq 0}$
10:     if i$=$max_iter then $\mathcal{D}\leftarrow\mathcal{D}\cup(\bm{X},Y,Z)$
11:     if $|\mathcal{D}|\geq N_{\text{init}}$ then take first $N_{\text{init}}$ and break
12:end for
13:return $f,q,\mathcal{D}_{0}=\mathcal{D}$ ($|\mathcal{D}|=N_{\text{init}}$)
@@ 3.2.1 Safety function $q$ and initial data $\mathcal{D}_{0}$
The first challenge is to ensure that that the simulated $q$ and $\mathcal{D}_{0}$ are both realistic and representative. In safe exploration problems [46, 55, 7], the initial data are usually provided by a domain expert at the centric area of a safe set and the algorithms gradually explore towards the safe set border. Following this principle, our generative algorithm samples initial data $\mathcal{D}_{0}$ in a pre-defined safe set, and we assume w.l.o.g. that the safe set lies at the center of $\mathcal{X}$, denoted by $\mathcal{C}\subseteq\mathcal{X}$. This approach ensures that $\mathcal{D}_{0}$ closely resembles the initial datasets encountered in the deployment stage.
@@ 3.2.2 Fourier feature functions - efficient and decoupled
The final step is to devise an efficient sampling strategy for the noisy GP values $y_{\phi,1:T_{\text{sim}}}$ and $z_{\phi,1:T_{\text{sim}}}$. This is not trivial because the observations are sampled iteratively, e.g. $\bm{x}_{t}=\phi(\mathcal{D}_{t-1})$, meaning $y_{\phi,1:t-1}$, $z_{\phi,1:t-1}$ are sampled before we know $\bm{x}_{t},...,\bm{x}_{T}$. One can make standard GP posterior sampling conditioned on preceding samples, e.g. $y_{t}\sim p(y(\bm{x}_{t})|\mathcal{D}_{t-1},k_{\theta},\sigma^{2})$. However, when performed naively, this posterior sampling approach is bound to the GP cubic complexity for each data point acquisition $y_{t}$, which is prohibitive. The runtime can be further reduced by using low-rank updates [41]. For specific objectives such as for our $\mathcal{H}$ or $\mathcal{S}_{\mathcal{H}}$, the GP posteriors can be then computed as intermediate results with minimal additional cost (see Eqs. 7 and 8). However, low-rank updates can lead to additional issues, as they make the computational graph unnecessarily complex for the loss function differentiation. To overcome these challenges, we propose a decoupled approach, which is scalable by itself and enables flexible integration of any objectives, e.g. potentially cheap approximated objectives, for future work.
The core is a decoupled function sampling technique [35, 53], where each GP function is approximated by a linear combination of Fourier features. The function value can be evaluated later at any $\bm{x}\in\mathcal{X}$ in linear time (line 8-9 of Algorithm 2). One limitation, however, is that the kernels $k_{\theta},k_{q}$ need to have Fourier transforms (e.g. stationary kernels, see Bochner’s theorem in [37]). For $q\sim\mathcal{GP}(\mu_{q},k_{q})$, one may sample $q_{\text{raw}}\sim\mathcal{GP}(0,k_{q})$ and set $q=\mu_{q}+q_{\text{raw}}$. Note that a sampled function does not necessarily average to zero on $\mathcal{X}$, but we may compute a domain specific average analytically with usually negligible complexity (details in Section D.1). This is preferred because we consider normalized deployment problems, and such an optional mean shift improves the performance.